# Topological photon pumping in quantum optical systems

### v2

@@ Abstract
We establish the concept of topological pumping in one-dimensional systems with long-range interactions and apply it to the transport of a photon in quantum optical systems. In our theoretical investigation, we introduce an extended version of the Rice-Mele model with all-to-all exchange interactions. By analyzing its properties, we identify the general conditions for topological pumping and demonstrate the topologically protected and dispersionless transport of a photon on a one-dimensional emitter chain. As concrete examples, we investigate three different popular quantum optics platforms, namely Rydberg atom lattices, dense lattices of atoms excited to low-lying electronic states, and atoms coupled to waveguides, using experimentally relevant parameters. We observe that despite the long-ranged character of the dipole-dipole interactions, topological pumping facilitates the transport of a photon with a fidelity per cycle which can reach 99.9%. Moreover, we find that the photon pumping process remains topologically protected against local disorder in the coupling rates.
@@ 1 Introduction
The precise control of the transport of a quantum state between different physical locations is a key ingredient for quantum information processing. Such transport can be for example implemented by moving atoms in optical tweezers [1, 2, 3, 4]. However, physically displacing quantum emitters over large distances is a delicate process and, ideally, such transport should be resilient to external disturbances. In the 80’s, Thouless [5] introduced the concept of topological charge pumps as a counter-intuitive, but robust transport mechanism founded on the principles of so-called topological protection [6]. Initially, the concept of topological charge pumps or topological pumping was formulated for solid-state systems involving electron transport, but over time it has been adapted to various systems, with ‘charge’ denoting a generic particle or excitation. In topological pumping, a center of mass displacement is introduced by adiabatically varying the system’s parameters in a way that allows controlled transitions between two topologically distinct phases which are characterized by topological invariants, such as the Chern number. When realized in a physical system, these topological invariants determine many of its physical properties, which in turn show remarkable robustness against local perturbations such as disorder [7, 8]. This robustness arises from mathematical symmetries - such as sublattice (chiral) or time-reversal symmetry - and has already been exploited for the fabrication of more robust devices [9, 10] and the establishment of physical standards [11].
@@
Only over the last decade, engineered systems such as waveguide arrays [12, 13, 14, 15], electronic systems [16, 17] and cold atoms systems [18, 19, 20] offered sufficient dynamical control of the system’s parameter to experimentally realize topological charge pumps. The displacement arises when the parameters are varied adiabatically around a topological degeneracy point. When an energy band is fully or uniformly occupied, the average displacement is given by an integer number of lattice sites [18, 19]. For inhomogeneous occupation of the bands, the displacement per cycle can be adjusted [20, 21]. In all these systems, the Hamiltonians are either non-interacting or display short range interactions (nearest neighbor hopping on a lattice, for example) for which the topological characterization and classification is relatively well understood [22, 23, 24, 25]. Recent advancements have extended this understanding to more complex non-Hermitian Hamiltonians [26, 27, 28, 29] and Hamiltonians with on-site interactions [30, 31, 32]. Conversely, very little is known about the fate of topological properties in systems with long-range interactions. Recent works indicate that the effect of such long-ranged character varies from system to system [33, 34, 35, 36]. Strikingly, it has been recently shown that some of the topological properties survive even in the presence of infinite range interactions, which can be found in, e.g. cavity and waveguide QED setups [37, 38, 39].
@@
Understanding the interplay between topology and symmetry-breaking long-range interactions extends well beyond the fundamental interest in condensed matter physics. This affects particularly the field of quantum optics, where systems inherently exhibit long-range dipole-dipole interactions. These systems constitute nowadays the vanguard of platforms for quantum simulation and quantum information processing with quantum emitters [40, 41], which can be realized with cold atoms [42, 43, 44], molecules [45, 46, 47] or resonators [48, 10, 49] placed in different one- two- and three-dimensional geometries. The coupling between these few-level emitters to a common environment can occur naturally, like in the case of the free space radiation field [50, 51, 52, 53, 54] or be engineered via optical and microwave cavities [55, 56], tapered optical waveguides [57, 58], and photonic crystals [59, 60, 43, 61], among others (see Fig. 1a). It is therefore an open question whether the promises of topology and in particular topological pumping can be applied in such systems.
@@
In this work, we investigate and obtain the conditions necessary to realize topological pumping in systems with long-range exchange interactions (Sec. 3). Moreover, we apply these results to demonstrate for the first time the topological pumping of a photon embedded in a one dimensional chain of emitters, i.e., how its position can be moved in a controlled way, while also being topologically protected against local disorder. Crucially, we take care that this transport is also dispersionless, such that not only the center of the photon wave packet is transported, but also its shape is conserved over several pump cycles, allowing for a high fidelity reaching 99.9% per cycle. We study this transport mechanism on three explicit platforms (Sec. 4), namely Rydberg atoms, atoms excited to low-lying electronic states, and atoms coupled to nanophotonic waveguides, which represent a relevant portion of the currently implemented experiments within the quantum optics community. Finally, we show that the topological transport mechanism is robust against dissipation and disorder (Sec. 5).
@@ 2 Topological pumping in one dimension
where |$\ket{u_{m}(t)}$ are the instantaneous eigenstates of the Hamiltonian. This geometric phase can cause a displacement of the average position of the state in real space after the cycle given by $\Delta x=\langle\hat{x}\rangle_{T}-\langle\hat{x}\rangle_{0}$, where $\langle\hat{x}\rangle_{t}\equiv\bra{\Psi_{t}}\hat{x}\ket{\Psi_{t}}$, which constitutes the so-called charge pumping. Note that, in general, also a dynamical phase may be present, which in turn may lead to a further displacement of the average position. In this paper, however, we restrict ourselves to situations where the contribution to the displacement of this dynamical phase can be neglected.
In one-dimensional periodic systems like the ones we will consider in this paper, it is convenient to work in quasi-momentum space. The eigenstates of the Hamiltonian are Bloch waves of the form |$\ket{\psi_{nk}(x,t)}=e^{\mathrm{i}kx}\ket{x}\otimes\ket{u_{nk}(t)}$, parametrized by their band index $n$ and the crystal quasi-momentum $\hbar k$. Here, |$\ket{x}$ is the position eigenstate ($\hat{x}\ket{x}=x\ket{x}$), and |$\ket{u_{nk}(t)}$ the cell-periodic part of the wave function. We consider the state of the dynamics at any time to be a superposition of these Bloch waves, i.e., |$\ket{\Psi_{t}(x)}=\sum_{k}f(k)\ket{\psi_{nk}(x,t)}$, where $f(k)$ is the quasi-momentum distribution amplitude, normalized such that $\int_{\mathrm{FBZ}}|f(k)|^{2}\mathrm{d}k=2\pi/a$ As detailed in Appendix A, one finds, for a general $f(k)$ distribution, that the displacement of the average position after a cycle reads
@@
$\Omega^{n}_{tk}=-2\text{Im}\{\langle\partial_{t}u_{nk}(t)|\partial_{k}u_{nk}(t)\rangle\},$ (3)
is the so-called Berry curvature of the $n$-th band [64, 63]. Note that, with a few notable examples [20, 65, 66], typically the literature of topological pumping considers a filled or uniformly populated band, where $f(k)$ is a constant function. In this case, the charge displacement is quantized in the form of $\Delta x=aC_{n}$, where $C_{n}$ is an integer called the Chern number of the band, defined as the integral of the Berry curvature over the two-dimensional torus formed by the FBZ and the periodic cycle in time. Finally, let us remark that for the displacement to be nonzero, a topological degeneracy point, i.e. the point in parameter space where the topological invariant changes, must be placed inside this torus [67].
Figure 2: Topological pumping in the Rice-Mele model. a: Sketch of the Rice-Mele model. $N/2$ cells formed by two sublattices, $A$ and $B$ (blue and red, respectively) are placed periodically with nearest neighbor distance $a$. Fermions can hop between sublattices with rates $J^{\prime}_{1}$ and $J_{1}$, and the sublattices are energetically offset by $2\hbar\Delta$. b: Adiabatically varying the parameters $\delta J=J_{1}^{\prime}-J_{1}$ and $\Delta$ in time in a closed cycle around the topological singularity point [orange dot at $(\Delta,\delta J)=(0,0)$], one can achieve topologically protected displacement of the fermions by an amount $\Delta x=a$ per cycle, as shown in panel c.
@@ 2.1 Rice-Mele model
The emblematic model Hamiltonian for charge pumping in one dimension is the so-called Rice-Mele model. Here, a one dimensional chain with $N$ sites is formed by two sublattices, $A$ and $B$. Fermions can hop between sublattices in the same cell and between neighboring cells with hopping rates $J^{\prime}_{1}$ and $J_{1}$, respectively (see Fig. 2a). In addition, an energy offset $2\hbar\Delta$ is introduced between the two sublattices. This system is described by the Hamiltonian
@@
In the Rice-Mele model where both $J_{1}$ and $J_{1}^{\prime}$ have the same sign, charge (fermion) pumping can be implemented when the parameters are varied in a cycle around the topological degeneracy point at $(\Delta,\delta J=J_{1}^{\prime}-J_{1})=(0,0)$ as sketched in Fig. 2b. Such a pumping scheme is robust against perturbations which deform the path, as long as it encircles the singularity. As sketched in Fig. 2c, when the band is filled, i.e. $f(k)$ constant in $k$, the center of mass of the wave packet is displaced after one cycle exactly by $\Delta x=a$, with $a$ being the distance between two neighboring unit cells. This quantized displacement is consistent with $\Delta x=aC_{n}$ since the Chern number of the lowest band is indeed $C_{n}=+1$ [67].
Figure 3: Topological pumping and dispersion. Left column: Time-integrated Berry curvature $\mathcal{W}^{n}_{k}$ and quasi-momentum distribution density $|f(k)|^{2}$. Right column: Spatial probability density of the state, $|\Psi_{t}(x)|^{2}\equiv{}\langle\Psi_{t}(x)|\Psi_{t}(x)\rangle$, at times $t=0$ and $t=5T$ corresponding to the quasi-momentum distribution density and time-integrated Berry curvature in the corresponding left panel. The vertical points and solid lines indicate the center of mass of the wave packet at $t=0$ and $5T$, respectively. a and b: In standard topological pumping, where the band is uniformly filled, the non-flat shape of ${\cal W}_{k}^{n}$ across the FBZ leads to dispersion of the wave packet in real space. Dispersionless topological pumping can be achieved either by making ${\cal W}_{k}^{n}$ constant (c and d), or by constraining $|f(k)|^{2}$ to a domain in momentum space where ${\cal W}_{k}^{n}$ is approximately constant (e and f).
@@ 2.2 Topological pumping and dispersion
Let us start by analysing the dispersion of the initial wave packet in real space |$\ket{\Psi_{0}(x)}=\sum_{k}f(k)\ket{\psi_{nk}(x,0)}$. After a full period, the state becomes |$\ket{\Psi_{T}(x)}=\sum_{k}f(k)e^{\mathrm{i}\gamma(k)}\ket{\psi_{nk}(x,0)}$, where
@@
When the initial state is such that a band is completely or uniformly filled [6, 66] with $\left|f(k)\right|^{2}$ constant for all $k$ (Fig. 3a), the charge displacement is quantized in the form of $\Delta x=aC_{n}$. However, in general ${\cal W}_{k}^{n}$ is not flat, which leads to a quick dispersion of an initially localized wave packet in real space (see Fig. 3b). Conversely, as discussed above, if ${\cal W}_{k}^{n}$ is constant throughout the full FBZ (see Fig. 3c), quantized and dispersionless pumping will be achieved independently of the specific shape of the wave packet in $k$ space (see Fig. 3d). This is, however, hard to achieve in practice in a real physical system, and even harder for a long-ranged Hamiltonian, which develops sharp structures in the Berry curvature.
In this paper, we resort to an intermediate approach, typically called geometric pumping [20]. For each system, we identify a path that leads to an time-integrated Berry curvature ${\cal W}_{k}^{n}$ that is constant on a finite region of the FBZ. We consider then a localized quasi-momentum distribution $f(k)$ such that $f(k)\neq 0$ only in the region where ${\cal W}_{k}^{n}$ is constant (see Fig. 3e) [21]. This condition is much easier to fulfill than a fully flat ${\cal W}_{k}^{n}$ [66], and allows an approximately dispersionless transport of the wave packet in real space. Instead of obtaining a displacement $\Delta x$ that is quantized, its value will depend on the specific geometrical properties of the underlying Hamiltonian (e.g., short versus long-ranged) as well as on the choice of the pumping cycle, since this determines the shape and value of ${\cal W}_{k}^{n}$ (see Fig. 3f). By choosing different protocols one can then achieve dispersionless transport of the wave packet where $\Delta x$ is larger or smaller than the one achieved in topological pumping.
Finally, note that the cycle period $T$ needs to be slow enough to ensure adiabaticity. This is fulfilled when $T\gg\hbar/\Delta E$, where $\Delta E$ is the minimum energy gap between neighboring bands during the cycle. On the other hand, for the transport to be as dispersionless as possible we need to consider an additional constraint: The cycle needs to be fast enough such that the small non-linearities inevitably present in the Berry phase $\gamma(k)$ do not lead to the dispersion of the wave packet. Throughout the paper, we have found cycle periods that approximately optimize these two conditions, although this still leads to a small degree of dispersion.

### v3

@@ Abstract
We establish the concept of topological pumping in one-dimensional systems with long-range couplings and apply it to the transport of a photon in quantum optical systems. In our theoretical investigation, we introduce an extended version of the Rice-Mele model with all-to-all couplings. By analyzing its properties, we identify the general conditions for topological pumping and theoretically and numerically demonstrate topologically protected and dispersionless transport of a photon on a one-dimensional emitter chain. As concrete examples, we investigate three different popular quantum optics platforms, namely Rydberg atom lattices, dense lattices of atoms excited to low-lying electronic states, and atoms coupled to waveguides, using experimentally relevant parameters. We observe that despite the long-ranged character of the dipole-dipole interactions, topological pumping facilitates the transport of a photon with a fidelity per cycle which can reach 99.9%. Moreover, we find that the photon pumping process remains topologically protected against local disorder in the coupling parameters.
@@ 1 Introduction
The precise control of the transport of a quantum state between different physical locations is a key ingredient for quantum information processing. Such transport can be for example implemented by moving atoms in optical tweezers [1, 2, 3, 4]. However, physically displacing quantum emitters over large distances is a delicate process and, ideally, such transport should be resilient to external disturbances. In the 80’s, Thouless [5] introduced the concept of topological charge pumps as a counter-intuitive, but robust transport mechanism founded on the principles of so-called topological protection [6]. Initially, the concept of topological charge pumps or topological pumping was formulated for solid-state systems involving electron transport, but over time it has been adapted to various systems, with ‘charge’ denoting a generic particle or excitation. In topological pumping, a center of mass displacement is introduced by adiabatically varying the system’s parameters in a way that allows controlled transitions between two topologically distinct phases which are characterized by topological invariants, such as the Chern number. When realized in a physical system, these topological invariants determine many of its physical properties, which in turn show remarkable robustness against local perturbations such as disorder [7, 8]. This robustness arises from mathematical symmetries - such as inversion or time-reversal symmetry - and has already been exploited for the fabrication of more robust devices [9, 10] and the establishment of physical standards [11].
@@
Only over the last decade, engineered systems such as waveguide arrays [12, 13, 14, 15], electronic systems [16, 17] and cold atoms systems [18, 19, 20] offered sufficient dynamical control of the system’s parameters to experimentally realize topological charge pumps. The displacement arises when the parameters are varied adiabatically around a topological degeneracy point. When an energy band is fully or uniformly occupied, the average displacement is given by an integer number of lattice sites [18, 19]. For inhomogeneous occupation of the bands, the displacement per cycle can be adjusted [20, 21]. In all these systems, the Hamiltonians are either non-interacting or display short range couplings (nearest neighbor hopping on a lattice, for example) for which the topological characterization and classification is relatively well understood [22, 23, 24, 25]. Recent advancements have extended this understanding to more complex non-Hermitian Hamiltonians [26, 27, 28, 29] and Hamiltonians with on-site interactions [30, 31, 32]. Conversely, very little is known about the fate of topological properties in systems with long-range couplings. Recent works indicate that the effect of such long-ranged character varies from system to system [33, 34, 35, 36]. Strikingly, it has been recently shown that some of the topological properties survive even in the presence of infinite range couplings, which can be found in, e.g. cavity and waveguide QED setups [37, 38, 39].
@@
Understanding the interplay between topology and long-range couplings extends well beyond the fundamental interest in condensed matter physics. This affects particularly the field of quantum optics, where systems inherently exhibit long-range dipole-dipole interactions. These systems constitute nowadays the vanguard of platforms for quantum simulation and quantum information processing with quantum emitters [40, 41], which can be realized with cold atoms [42, 43, 44], molecules [45, 46, 47] or resonators [48, 10, 49] placed in different one- two- and three-dimensional geometries. The coupling between these few-level emitters to a common environment can occur naturally, like in the case of the free space radiation field [50, 51, 52, 53, 54] or be engineered via optical and microwave cavities [55, 56], tapered optical waveguides [57, 58], and photonic crystals [59, 60, 43, 61], among others (see Fig. 1a). It is therefore an open question whether the promises of topology and in particular topological pumping can be applied in such systems.
@@
In this work, we investigate and obtain the conditions necessary to realize topological pumping in systems with long-range couplings (Sec. 3). Moreover, we apply these results to theoretically demonstrate for the first time the topological pumping of a photon embedded in a one dimensional chain of emitters, i.e., how its position can be moved in a controlled way, while also being topologically protected against local disorder. Crucially, we take care that this transport is also dispersionless, such that not only the center of the photon wave packet is transported, but also its shape is conserved over several pump cycles, allowing for a high fidelity reaching 99.9% per cycle. We study this transport mechanism on three explicit platforms (Sec. 4), namely Rydberg atoms, atoms excited to low-lying electronic states, and atoms coupled to nanophotonic waveguides, which represent a relevant portion of the currently implemented experiments within the quantum optics community. Finally, we show that the topological transport mechanism is robust against dissipation and disorder (Sec. 5).
@@ 2 Topological pumping in one dimension
where |$\ket{u_{m}(t)}$ are the instantaneous eigenstates of the Hamiltonian. This geometric phase can cause a displacement of the average position of the state in real space after the cycle given by $\Delta x=\langle\hat{x}\rangle_{T}-\langle\hat{x}\rangle_{0}$, where $\langle\hat{x}\rangle_{t}\equiv\bra{\Psi_{t}}\hat{x}\ket{\Psi_{t}}$, which constitutes the so-called charge pumping.
In one-dimensional periodic systems like the ones we will consider in this paper, it is convenient to work in quasi-momentum space. The eigenstates of the Hamiltonian are Bloch waves of the form |$\ket{\psi_{nk}(x,t)}\propto e^{\mathrm{i}kx}\ket{x}\otimes\ket{u_{nk}(t)}$, parametrized by their band index $n$ and the crystal quasi-momentum $\hbar k$. Here, |$\ket{x}$ is the position eigenstate ($\hat{x}\ket{x}=x\ket{x}$), and |$\ket{u_{nk}(t)}$ the cell-periodic part of the wave function. We consider the state of the dynamics at any time to be a superposition of these Bloch waves, i.e., |$\ket{\Psi_{t}(x)}\propto\sum_{k}f(k)\ket{\psi_{nk}(x,t)}$, where $f(k)$ is the quasi-momentum distribution amplitude. As detailed in Appendix A, one finds, for a general $f(k)$ distribution, that the displacement of the average position after a cycle reads
@@
$\Omega^{n}_{tk}=-2\text{Im}\{\langle\partial_{t}u_{nk}(t)|\partial_{k}u_{nk}(t)\rangle\}$ (3)
is the so-called Berry curvature of the $n$-th band [64, 63]. With a few notable exceptions [20, 65, 66], typically the literature of topological pumping considers a filled band or a single-particle wave packet that is delocalized in momentum space, such that $f(k)$ is a constant function. In this case, the charge displacement is quantized in the form of $\Delta x=aC_{n}$, where $C_{n}$ is an integer called the Chern number of the band, defined as the integral of the Berry curvature over the two-dimensional torus formed by the FBZ and the periodic cycle in time. Let us remark that for the displacement to be nonzero, a topological degeneracy point, i.e. the point in parameter space where the topological invariant changes, must be placed inside this torus [67]. Finally, note that, in general, also a dynamical phase may be present, which is proportional to the usual group velocity. This in turn may lead to a further displacement of the average position, given by the convolution of the group velocity with the quasi-momentum distribution, integrated over the FBZ. In this paper, however, we are interested in the anomalous displacement arising from the topological properties and an adiabatic change of the system. We keep this effect dominant by restricting ourselves to symmetric wave packets placed in the center of the FBZ of a time-reversal symmetric Hamiltonian, which leads to a zero contribution to the displacement of the dynamical phase.
Figure 2: Topological pumping in the Rice-Mele model. a: Sketch of the Rice-Mele model. $N/2$ cells formed by two sublattices, $A$ and $B$ (blue and red, respectively) are placed periodically with nearest neighbor distance $a$. Fermions can hop between sublattices (with hopping parameters $J^{\prime}_{1}$ and $J_{1}$), and the sublattices are energetically offset by $2\hbar\Delta$. b: Adiabatically varying the parameters $\delta J=J_{1}^{\prime}-J_{1}$ and $\Delta$ in time in a closed cycle around the topological singularity point [orange dot at $(\Delta,\delta J)=(0,0)$], one can achieve topologically protected displacement of the fermions by an amount $\Delta x=a$ per cycle, as shown in panel c.
@@ 2.1 Rice-Mele model
The emblematic model Hamiltonian for charge pumping in one dimension is the so-called Rice-Mele model. Here, a one dimensional chain with $N$ sites is formed by two sublattices, $A$ and $B$. Fermions can hop between sublattices in the same cell and between neighboring cells with hopping parameters $J^{\prime}_{1}$ and $J_{1}$, respectively (see Fig. 2a). In addition, an energy offset $2\hbar\Delta$ is introduced between the two sublattices. This system is described by the Hamiltonian
@@
In the Rice-Mele model where both $J_{1}$ and $J_{1}^{\prime}$ have the same sign, charge (fermion) pumping can be implemented when the parameters are varied in a cycle around the topological degeneracy point at $(\Delta,\delta J=J_{1}^{\prime}-J_{1})=(0,0)$ as sketched in Fig. 2b. Such a pumping scheme is robust against perturbations which deform the path, as long as it encircles the singularity. As sketched in Fig. 2c, when $f(k)$ constant in $k$, the center of mass of the wave packet is displaced after one cycle exactly by $\Delta x=a$, with $a$ being the distance between two neighboring unit cells. This quantized displacement is consistent with $\Delta x=aC_{n}$ since the Chern number of the lowest band is indeed $C_{n}=+1$ [67].
@@ 2.2 Topological pumping and dispersion
Let us start by analysing the dispersion of the initial wave packet in real space |$\ket{\Psi_{0}(x)}\propto\sum_{k}f(k)\ket{\psi_{nk}(x,0)}$. After a full period, the state becomes |$\ket{\Psi_{T}(x)}\propto\sum_{k}f(k)e^{\mathrm{i}\gamma(k)}\ket{\psi_{nk}(x,0)}$, where
@@
When the initial state is such that $\left|f(k)\right|^{2}$ constant for all $k$ (Fig. 3a) [6, 66], the charge displacement is quantized in the form of $\Delta x=aC_{n}$. However, in general ${\cal W}_{k}^{n}$ is not flat, which leads to a quick dispersion of an initially localized wave packet in real space (see Fig. 3b). Conversely, as discussed above, if ${\cal W}_{k}^{n}$ is constant throughout the full FBZ (see Fig. 3c), quantized and dispersionless pumping will be achieved independently of the specific shape of the wave packet in $k$ space (see Fig. 3d). This is, however, hard to achieve in practice in a real physical system, and even harder for a long-ranged Hamiltonian, which develops sharp structures in the Berry curvature.
In this paper, we resort to an intermediate approach, typically called geometric pumping [20]. For each system, we identify a path that leads to a time-integrated Berry curvature ${\cal W}_{k}^{n}$ that is constant on a finite region of the FBZ. We consider then a localized quasi-momentum distribution $f(k)$ such that $f(k)\neq 0$ only in the region where ${\cal W}_{k}^{n}$ is constant (see Fig. 3e) [21]. This condition is much easier to fulfill than a fully flat ${\cal W}_{k}^{n}$ [66], and allows an approximately dispersionless transport of the wave packet in real space. Instead of obtaining a displacement $\Delta x$ that is quantized, its value will depend on the specific geometrical properties of the underlying Hamiltonian (e.g., short versus long-ranged) as well as on the choice of the pumping cycle, since this determines the shape and value of ${\cal W}_{k}^{n}$ (see Fig. 3f). By choosing different protocols one can then achieve dispersionless transport of the wave packet where $\Delta x$ is larger or smaller than the one achieved in topological pumping.
Figure 3: Topological pumping and dispersion. Left column: Time-integrated Berry curvature $\mathcal{W}^{n}_{k}$ and quasi-momentum distribution density $|f(k)|^{2}$. Right column: Spatial probability density of the state, $|\Psi_{t}(x)|^{2}\equiv{}\langle\Psi_{t}(x)|\Psi_{t}(x)\rangle$, at times $t=0$ and $t=5T$, obtained from the numerical simulation of the Schrödinger equation under a time-dependent Hamiltonian of the form (4), with the quasi-momentum distribution density and time-integrated Berry curvature shown in the corresponding left panel. The vertical points and solid lines indicate the center of mass of the wave packet at $t=0$ and $5T$, respectively. a and b: In standard topological pumping, where the band is uniformly filled, the non-flat shape of ${\cal W}_{k}^{n}$ across the FBZ leads to dispersion of the wave packet in real space. Dispersionless topological pumping can be achieved either by making ${\cal W}_{k}^{n}$ constant (c and d), or by constraining $|f(k)|^{2}$ to a domain in momentum space where ${\cal W}_{k}^{n}$ is approximately constant (e and f).